# Universally Optimal Periodic Configurations in the Plane

### v3

@@ 3.2 Interpolation Nodes
[…]
@@
$\tilde{\tau}_{m^{2}}$ =$=\left\{\left(\cos(\frac{\pi k_{1}}{m}),\cos(\frac{\pi k_{2}}{m})\right)\mid 0\leq 3k_{2}\leq k_{1}\leq m,k_{1}\equiv k_{2}(\bmod 2)\right\}$ (46)
$\tilde{\tau}_{2m^{2}}$ =$=\left\{\left(\cos(\frac{\pi k_{1}}{m}),\cos(\frac{\pi k_{2}}{m})\right)\mid 0\leq k_{1},k_{2}\leq m,k_{1}\equiv k_{2}(\bmod 2)\right\}$ (47)
$\tilde{\tau}_{3m^{2}}$ =$=\left\{\left(\cos(\frac{\pi k_{1}}{m}),\cos(\frac{\pi k_{2}}{3m})\right)\mid 0\leq k_{2}\leq k_{1}\leq m,k_{1}\equiv k_{2}(\bmod 2)\right\}$ (48)
$\tilde{\tau}_{6m^{2}}$ =$=\left\{\left(\cos(\frac{\pi k_{1}}{m}),\cos(\frac{\pi k_{2}}{3m})\right)\mid 0\leq 3k_{1},k_{2}\leq 3m,k_{1}\equiv k_{2}(\bmod 2)\right\}.$ (49)
@@ 3.3 Interpolation Problem for $\omega_{m^{2}}$
Figure 4: $\tilde{g}_{a}$ must stay below $\tilde{F}_{a}$ on $\tilde{\Delta}_{A_{2}}$ with equality at the corner point $(-1,1)$.
@@ 3.4 Interpolation Problem for $\omega^{*}_{2m^{2}}$
Figure 7: After applying the linear programming framework, it suffices to find for each $a>0$ an interpolant $\tilde{g}_{a}\in\mathcal{C}_{8}$ such that $\tilde{g}_{a}\leq\tilde{F}_{a}$ on $[-1,1]$ with equality at the 4 points shown.
@@ 3.5 Interpolation Problems for $\omega^{*}_{3m^{2}}$
The universal optimality of $\omega^{*}_{3}$ (cf. [33] and [16]) follows from Lemma 15, which is used to show a global minimum of $F_{a,A_{2}}$ occurs at $(1/2,\sqrt{3}/6)$ for all $a>0$. This point, $(1/2,\sqrt{3}/6)$, is also the only difference $x-y$ (up to $S_{A_{2}}$ action) for $x,y\in\omega^{*}_{3}$. Thus for an arbitrary 3-point configuration $\omega_{3}$, we have
@@ 3.6 Interpolation Problems for $\omega^{*}_{6m^{2}}$
Figure 10: $\tilde{g}_{a}$ must stay below $\tilde{F}_{a}$ on $[-1,1]^{2}$ with equality at the three points shown
@@ 4 $A_{2}$-universal optimality of $\omega^{*}_{4}$
To prove $\omega^{*}_{4}$ is $A_{2}$-universally optimal, it remains to show for each $a>0$ that there are $c_{0},c_{1}\in\mathbb{R}$ with $c_{1}\geq 0$ such that the resulting interpolant $\tilde{g}_{a}(t_{1},t_{2}):=c_{0}+c_{1}P_{v^{\prime}}=c_{0}+c_{1}/3(-1+t_{2}(t_{1}+t_{2}))$ satisfies $\tilde{g}_{a}\leq\tilde{F}_{a}$ on $\tilde{\Delta}_{A_{2}}$ with equality at $(-1,1)$ or, equivalently, finding such an interpolant of the form
$\ \tilde{g}_{a}(t_{1},t_{2}):=\tilde{F}_{a}(-1,1)+b_{1}t_{2}(t_{1}+t_{2})$ (50)
@@ 4.1 Constructing magic $\tilde{g}_{a}$
We abuse notation here and use $\tilde{F},\tilde{h}$ to refer to the one variable functions in $t_{2}$ obtained by fixing $t_{1}=-1$. By assumption on the form of $\tilde{h}$, Lemma 15, and the two assumed inequalities, we have $\tilde{F}(1/2)\geq\tilde{h}(1/2)$, $\tilde{F}^{\prime}(1/2)=\tilde{h}^{\prime}(1/2)=0$, $\tilde{F}(1)=\tilde{h}(1)$, and $\tilde{F}^{\prime}(1)\leq\tilde{h}^{\prime}(1)$. It follows that there are exists some point in $[1/2,1]$ at which $\tilde{F}^{\prime\prime}\leq\tilde{h}^{\prime\prime}$. Let $t_{2}^{\prime}\leq t_{2}^{\prime\prime}$ be such that $t_{2}^{\prime}$ and $t_{2}^{\prime\prime}$ respectively are the minimal and maximal points in $[1/2,1]$ at which $\tilde{F}^{\prime\prime}\leq\tilde{h}^{\prime\prime}$. For $t_{2}\geq t_{2}^{\prime\prime}$ we have $\tilde{F}^{\prime\prime}\geq\tilde{h}^{\prime\prime}$ with equality only at $t_{2}^{\prime\prime}$ since $(\tilde{F}-\tilde{h}){{}^{\prime\prime}}$ is strictly convex (recall $\tilde{F}^{(4)}>0$). Thus, we get $\tilde{F}\geq\tilde{h}$ by bounding $\tilde{F}-\tilde{h}$ below with a tangent line of $\tilde{F}-\tilde{h}$ at $1$ and equality holds only if $t_{2}=1$. Similarly, for $t_{2}\leq t_{2}^{\prime}$, we get the desired inequality with tangent approximation from $\frac{1}{2}$. For $t_{2}\in[t_{2}^{\prime},t_{2}^{\prime\prime}]$, we note that $\tilde{F}^{\prime\prime}=\tilde{h}^{\prime\prime}$ at the endpoints of the interval. Again using the strict convexity of $(\tilde{F}-\tilde{h}){{}^{\prime\prime}}$, we obtain $(\tilde{F}-\tilde{h})^{\prime\prime}\leq 0$ for the whole interval. Thus, $\tilde{F}(t_{2})\geq\tilde{h}(t_{2})$ by bounding the difference below with its secant line (since we’ve already established $\tilde{F}\geq\tilde{h}$ at the endpoints $t_{2}^{\prime},t_{2}^{\prime\prime}$), and equality can only hold at $t_{2}^{\prime\prime}$ if $t_{2}^{\prime\prime}=1$. ∎
@@ 4.1.1 Small $a$
$\frac{\partial\tilde{F}}{\partial t_{2}}(-1,1),4(\tilde{F}(-1,1)-\tilde{F}(-1,1/2))\leq 2\frac{\partial\tilde{F}}{\partial t_{1}}(-1,1/2)\leq\frac{\partial^{2}(\tilde{F}-\tilde{g})}{\partial t_{1}\partial t_{2}}(-1,1/2).$
@@
Applying Lemma 17 and the previous two inequalities to $\tilde{g}$, we obtain $\tilde{F}\geq\tilde{g}$ for all points $(-1,t_{2})$ with $t_{2}\in[1/2,1]$ with equality only at $(-1,1)$, and since $\tilde{F}-\tilde{g}$ is convex in $t_{1}$, it remains to show that $\frac{\partial(\tilde{F}-\tilde{g})}{\partial t_{1}}\geq 0$ for all points of the form $(-1,t_{2})$, $t_{2}\in[1/2,1]$ (recall the picture of $\tilde{\Delta}_{A_{2}}$, Figure 4). By Lemma 16, $\frac{\partial(\tilde{F}-\tilde{g})}{\partial t_{1}}$ is convex in $t_{2}$ in $\tilde{\Delta}_{A_{2}}$, so we just need to show
@@ 4.1.2 Large $a$
We have $\tilde{F}\geq\tilde{g}$ on the set $\{(-1,t_{2}):t_{2}\in[1/2,1]\}\cup\{(t_{1},1):t_{1}\in[-1,1]\}$ with equality only at $(-1,1)$.
@@
$\frac{\partial\tilde{F}}{\partial t_{1}}(-1,1)>\frac{\partial\tilde{F}}{\partial t_{2}}(-1,1)$ (56)
It also holds for $210<a<$ 21 as an immediate consequence of Lemma 18. We next show for $a>21$ that $(\tilde{F}-\tilde{g})(-1,1/2)>0$, and so using the definition
@@
for this range of $a$, we may apply Lemma 17 to obtain $\tilde{F}\geq\tilde{g}$ on $\{(-1,t_{2}):t_{2}\in[1/2,1]\}$ with equality only at $(-1,1)$. Now for the other segment, we simply apply (56), our definition of $b_{1}$, and the convexity of $\tilde{F}-\tilde{g}$ in $t_{1}$. ∎
Next, we show that $\tilde{F}(t_{1},t_{2})\geq\tilde{g}(t_{1},t_{2})$ in $\tilde{\Delta}_{A_{2}}$ if $t_{1}\leq\cos(2\pi\frac{\sqrt{3}}{4})$ with equality only at $(-1,1)$, and in fact we’ll show the stronger claim that $\tilde{F}\geq\tilde{g}$ on all of $A:=[-1,\cos(2\pi\frac{\sqrt{3}}{4})]\times[1/2,1]$ (see Figure 11) with equality only at $(-1,1)$.
@@
hold on $\partial R$, then $\tilde{F}>\tilde{g}$ on $R$ for all $a\geq a^{\prime}$. Further, if for all $a\geq a^{\prime}$, $\tilde{F}\geq\tilde{g}$ on some $R^{\prime}\subseteq\{(t_{1},t_{2})\in\partial R:t_{1}=c\text{ or }t_{2}=d\}$ and the inequalities (62) hold on $\partial R\backslash R^{\prime}$, then $\tilde{F}\geq\tilde{g}$ on $R$ for all $a\geq a^{\prime}$, again with equality only possible if $t_{1}=c$ or $t_{2}=d$ and $(t_{1},t_{2})\in R^{\prime}$.
@@
The inequality $\tilde{F}\geq\tilde{g}$ holds on $A=[-1,\cos(2\pi\frac{\sqrt{3}}{4})]\times[\frac{1}{2},1]$ with equality only at $(-1,1)$.
@@
We partition $[-1,\cos(2\pi\frac{\sqrt{3}}{4})]\times[\frac{1}{2},1]$ into three subrectangles $R_{k}:=[-1,\cos(2\pi\frac{\sqrt{3}}{4})]\times[d_{k-1},d_{k}]$, $k=1,2,3$ where $d_{0}=1/2$, $d_{1}=3/5$, $d_{2}=7/10$, and $d_{3}=1$ and aim to verify the inequality $\tilde{F}\geq\tilde{g}_{-1,d_{k}}$ on each $R_{k}$ using Lemma 20, with $\tilde{g}_{-1,d_{k}}$ as in (57), $\tilde{F}_{T}$ as in (59), $m_{1}=4$, and $a^{\prime}=21$. The specific formulas for each $g_{-1,d_{k}}^{*}$ are given in the appendix section D.3. We begin by verifying inequalities (62) for $g_{-1,-1}^{*}$ on the line segments of $R_{1}$ with $t_{1}=\cos(2\pi\sqrt{3}/4)$ or $t_{2}=7/10$, which combined with Lemma 20 implies $\tilde{F}\geq\tilde{g}$ on $R_{1}$. Now having established $\tilde{F}\geq\tilde{g}$ on the top side of $R_{2}$, we only need establish inequalities (62) for $g_{-1,-1}^{*}$ on the line segments of $R_{2}$ where $t_{1}=\cos(2\pi\sqrt{3}/4)$ or $t_{2}=3/5$ to get $\tilde{F}\geq\tilde{g}$ on all $R_{2}$. In the same fashion, showing inequalities (62) on the sides of $R_{3}$ where $t_{1}=\cos(2\pi\sqrt{3}/4)$ or $t_{2}=3/5$ completes the proof by yielding $\tilde{F}\geq\tilde{g}$ on $R_{3}$. The verification of these inequalities is carried out in the appendix by reducing them to inequalities of the form $h_{2}(t)-h_{1}(t)>0$ on an interval $(\alpha,\beta)$ where $h_{1}$ and $h_{2}$ are increasing functions and choose $\delta:=(\beta-\alpha)/n$ with sufficiently large that we may rigorously verify the inequalities
@@
Finally, we show that $\tilde{F}-\tilde{g}$ increases in $t_{1}$ for every point in $\tilde{\Delta}_{A_{2}}$ with $t_{1}\geq\cos(2\pi\frac{\sqrt{3}}{4})$, thus completing the proof for the $a>21$ case and yielding $\tilde{F}\geq\tilde{g}$ on all $\tilde{\Delta}_{A_{2}}$ with equality only at $(-1,1)$.
@@
[…]
@@ 5 $L$-universal optimality of $\omega^{*}_{6}$
We require that $\tilde{g}$ and $\tilde{F}$ agree at the three points in $\tilde{\tau}_{6}$ and remark that the condition $\tilde{g}\leq\tilde{F}$ further requires $\partial\tilde{g}/\partial t_{2}=\partial\tilde{F}/\partial t_{2}$ at the points $(-1,1/2)$ and $(1,-1/2)$ giving a total of 5 linearly independent conditions on $\mathcal{P}_{1}(t_{1})\times\mathcal{P}_{2}(t_{2})$.
@@
and $H_{\{-\frac{1}{2},-\frac{1}{2}\}}(\tilde{f_{2}})(t_{2})=\tilde{f_{2}}(-\frac{1}{2})+\tilde{f_{2}}^{\prime}(-\frac{1}{2})(t_{2}+\frac{1}{2})^{2}$. Since $T_{1}(t)=t$ and $T_{2}(t)=2t^{2}-1$, it easily follows that $\tilde{g}$ is CPSD if and only if $b_{i,j}\geq 0$ for $(i,j)\neq 0$. From (65), it follows that $b_{1,2}=-\frac{1}{2}\tilde{f_{1}}(-1)\tilde{f_{2}}[-1,\frac{1}{2},\frac{1}{2}]+c$. Observing that $q\geq 0$ on $[-1,1]^{2}$, we choose $c=\frac{1}{2}\tilde{f_{1}}(-1)\tilde{f_{2}}[-1,\frac{1}{2},\frac{1}{2}]$ as small as possible in which case $b_{1,2}=0$.
@@
The next lemma which will be used to prove $a_{1,0}>0$ as well as being a first step in establishing that $\tilde{g}\leq\tilde{F}$ on $[-1,1]^{2}$.
@@
5.1 $\tilde{F}\geq\tilde{g}$ on $[-1,1]\times([-1,-1/2]\cup[1/2,1])$
The following lemma, proved in the appendix, establishes several necessary inequality conditions for $\tilde{g}\leq\tilde{F}$. We next show the inequality holds on $[-1,1]\times([-1,-1/2]\cup[1/2,1])$.

### v4

@@ 3.2 Interpolation Nodes
$\tau_{\kappa m^{2}}$ $=\omega_{\kappa m^{2}}\cap([0,1/2]\times[0,\sqrt{3}/2]),\hskip 28.45274pt\kappa=2,6$ (45)
@@
$\tilde{\tau}_{m^{2}}$ $=\left\{\left(\cos(\frac{\pi k_{1}}{m}),\cos(\frac{\pi k_{2}}{m})\right)\mid 0\leq 3k_{2}\leq k_{1}\leq m,k_{1}\equiv k_{2}(\bmod 2)\right\}$ (46)
$\tilde{\tau}_{2m^{2}}$ $=\left\{\left(\cos(\frac{\pi k_{1}}{m}),\cos(\frac{\pi k_{2}}{m})\right)\mid 0\leq k_{1},k_{2}\leq m,k_{1}\equiv k_{2}(\bmod 2)\right\}$ (47)
$\tilde{\tau}_{3m^{2}}$ $=\left\{\left(\cos(\frac{\pi k_{1}}{m}),\cos(\frac{\pi k_{2}}{3m})\right)\mid 0\leq k_{2}\leq k_{1}\leq m,k_{1}\equiv k_{2}(\bmod 2)\right\}$ (48)
$\tilde{\tau}_{6m^{2}}$ $=\left\{\left(\cos(\frac{\pi k_{1}}{m}),\cos(\frac{\pi k_{2}}{3m})\right)\mid 0\leq 3k_{1},k_{2}\leq 3m,k_{1}\equiv k_{2}(\bmod 2)\right\}.$ (49)
@@ 3.3 Interpolation Problem for $\omega_{m^{2}}$
Figure 4: $\tilde{g}_{a}$ must stay below $\tilde{F}_{a}$ on $\tilde{\Delta}_{A_{2}}$ with equality at the corner point (1,1)(-1,1).
@@ 3.4 Interpolation Problem for $\omega^{*}_{2m^{2}}$
Figure 7: After applying the linear programming framework, it suffices to find for each $a>0$ an interpolant $\tilde{g}_{a}\in\mathcal{C}_{8}$ such that $\tilde{g}_{a}\leq\tilde{F}_{a}$ on [1,1][-1,1] with equality at the 4 points shown.
@@ 3.5 Interpolation Problems for $\omega^{*}_{3m^{2}}$
The universal optimality of $\omega^{*}_{3}$ (cf. [22] and [11]) follows from Lemma 15, which is used to show a global minimum of $F_{a,A_{2}}$ occurs at $(1/2,\sqrt{3}/6)$ for all $a>0$. This point, $(1/2,\sqrt{3}/6)$, is also the only difference xyx-y (up to $S_{A_{2}}$ action) for $x,y\in\omega^{*}_{3}$. Thus for an arbitrary 3-point configuration $\omega_{3}$, we have
@@ 3.6 Interpolation Problems for $\omega^{*}_{6m^{2}}$
Figure 10: $\tilde{g}_{a}$ must stay below $\tilde{F}_{a}$ on $1,1]^{2}$ with equality at the three points shown
@@ 4 $A_{2}$-universal optimality of $\omega^{*}_{4}$
To prove $\omega^{*}_{4}$ is $A_{2}$-universally optimal, it remains to show for each $a>0$ that there are $c_{0},c_{1}\in\mathbb{R}$ with $c_{1}\geq 0$ such that the resulting interpolant $\tilde{g}_{a}(t_{1},t_{2}):=c_{0}+c_{1}P_{v^{\prime}}=c_{0}+c_{1}/3(-1+t_{2}(t_{1}+t_{2}))$ satisfies $\tilde{g}_{a}\leq\tilde{F}_{a}$ on $\tilde{\Delta}_{A_{2}}$ with equality at (1,1)(-1,1) or, equivalently, finding such an interpolant of the form
[…]
@@ 4.1 Constructing magic $\tilde{g}_{a}$
[…]
@@ 4.1.1 Small $a$
$\max\left\{\frac{\partial\tilde{F}}{\partial t_{2}}(-1,1),\hskip 7.11317pt4(\tilde{F}(-1,1)-\tilde{F}(-1,1/2))\right\}\leq 2\frac{\partial\tilde{F}}{\partial t_{1}}(-1,1/2)\leq\frac{\partial^{2}(\tilde{F}-\tilde{g})}{\partial t_{1}\partial t_{2}}(-1,1/2).$
@@
Applying Lemma 17 and the previous two inequalities to $\tilde{g}$, we obtain $\tilde{F}\geq\tilde{g}$ for all points $(-1,t_{2})$ with $t_{2}\in[1/2,1]$ with equality only at (1,1)(-1,1), and since $\tilde{F}-\tilde{g}$ is convex in $t_{1}$, it remains to show that $\frac{\partial(\tilde{F}-\tilde{g})}{\partial t_{1}}\geq 0$ for all points of the form $(-1,t_{2})$, $t_{2}\in[1/2,1]$ (recall the picture of $\tilde{\Delta}_{A_{2}}$, Figure 4). By Lemma 16, $\frac{\partial(\tilde{F}-\tilde{g})}{\partial t_{1}}$ is convex in $t_{2}$ in $\tilde{\Delta}_{A_{2}}$, so we just need to show
@@ 4.1.2 Large $a$
We have $\tilde{F}\geq\tilde{g}$ on the set $\{(-1,t_{2}):t_{2}\in[1/2,1]\}\cup\{(t_{1},1):t_{1}\in[-1,1]\}$ with equality only at (1,1)(-1,1).
@@
$\frac{\partial\tilde{F}}{\partial t_{1}}(-1,1)>\frac{\partial\tilde{F}}{\partial t_{2}}(-1,1).$ (56)
It also holds for $0<a<21$ as an immediate consequence of Lemma 18. We next show for $a>21$ that $(\tilde{F}-\tilde{g})(-1,1/2)>0$, and so using the definition
@@
for this range of $a$, we may apply Lemma 17 to obtain $\tilde{F}\geq\tilde{g}$ on $\{(-1,t_{2}):t_{2}\in[1/2,1]\}$ with equality only at (1,1)(-1,1). Now for the other segment, we simply apply (56), our definition of $b_{1}$, and the convexity of $\tilde{F}-\tilde{g}$ in $t_{1}$. ∎
Next, we show that $\tilde{F}(t_{1},t_{2})\geq\tilde{g}(t_{1},t_{2})$ in $\tilde{\Delta}_{A_{2}}$ if $t_{1}\leq\cos(2\pi\frac{\sqrt{3}}{4})$ with equality only at (1,1)(-1,1), and in fact we’ll show the stronger claim that $\tilde{F}\geq\tilde{g}$ on all of $A:=[-1,\cos(2\pi\frac{\sqrt{3}}{4})]\times[1/2,1]$ (see Figure 11) with equality only at (1,1)(-1,1).
@@
[…]
@@
The inequality $\tilde{F}\geq\tilde{g}$ holds on $A=[-1,\cos(2\pi\frac{\sqrt{3}}{4})]\times[\frac{1}{2},1]$ with equality only at (1,1)(-1,1).
@@
We partition $1,\cos(2\pi\frac{\sqrt{3}}{4})]\times[\frac{1}{2},1]$ into three subrectangles $R_{k}:=[-1,\cos(2\pi\frac{\sqrt{3}}{4})]\times[d_{k-1},d_{k}]$, $k=1,2,3$ where $d_{0}=1/2$, $d_{1}=3/5$, $d_{2}=7/10$, and $d_{3}=1$ and aim to verify the inequality $\tilde{F}\geq\tilde{g}_{-1,d_{k}}$ on each $R_{k}$ using Lemma 20, with $\tilde{g}_{-1,d_{k}}$ as in (57), $\tilde{F}_{T}$ as in (59), $m_{1}=4$, and $a^{\prime}=21$. The specific formulas for each $g_{-1,d_{k}}^{*}$ are given in the appendix section D.3. We begin by verifying inequalities (62) for $g_{-1,-1}^{*}$ on the line segments of $R_{1}$ with $t_{1}=\cos(2\pi\sqrt{3}/4)$ or $t_{2}=7/10$, which combined with Lemma 20 implies $\tilde{F}\geq\tilde{g}$ on $R_{1}$. Now having established $\tilde{F}\geq\tilde{g}$ on the top side of $R_{2}$, we only need establish inequalities (62) for $g_{-1,-1}^{*}$ on the line segments of $R_{2}$ where $t_{1}=\cos(2\pi\sqrt{3}/4)$ or $t_{2}=3/5$ to get $\tilde{F}\geq\tilde{g}$ on all $R_{2}$. In the same fashion, showing inequalities (62) on the sides of $R_{3}$ where $t_{1}=\cos(2\pi\sqrt{3}/4)$ or $t_{2}=3/5$ completes the proof by yielding $\tilde{F}\geq\tilde{g}$ on $R_{3}$. The verification of these inequalities is carried out in the appendix by reducing them to inequalities of the form $h_{2}(t)-h_{1}(t)>0$ on an interval $(\alpha,\beta)$ where $h_{1}$ and $h_{2}$ are increasing functions and choose $\delta:=(\beta-\alpha)/n$ with sufficiently large that we may rigorously verify the inequalities
@@
Finally, we show that $\tilde{F}-\tilde{g}$ increases in $t_{1}$ for every point in $\tilde{\Delta}_{A_{2}}$ with $t_{1}\geq\cos(2\pi\frac{\sqrt{3}}{4})$, thus completing the proof for the $a>21$ case and yielding $\tilde{F}\geq\tilde{g}$ on all $\tilde{\Delta}_{A_{2}}$ with equality only at (1,1)(-1,1).
@@
$\frac{\partial(\tilde{F}-\tilde{g})}{\partial t_{1}}\bigg|_{P}\geq 0,\hskip 28.45274pt\frac{\partial^{2}(\tilde{F}-\tilde{g})}{\partial t_{1}\partial t_{2}}\bigg|_{P}\geq 0$ (64)
@@ 5 $L$-universal optimality of $\omega^{*}_{6}$
We require that $\tilde{g}$ and $\tilde{F}$ agree at the three points in $\tilde{\tau}_{6}$ and remark that the condition $\tilde{g}\leq\tilde{F}$ further requires $\partial\tilde{g}/\partial t_{2}=\partial\tilde{F}/\partial t_{2}$ at the points (1,1/2)(-1,1/2) and (1,1/2)(1,-1/2) giving a total of 5 linearly independent conditions on $\mathcal{P}_{1}(t_{1})\times\mathcal{P}_{2}(t_{2})$.
@@
and $H_{\{-\frac{1}{2},-\frac{1}{2}\}}(\tilde{f_{2}})(t_{2})=\tilde{f_{2}}(-\frac{1}{2})+\tilde{f_{2}}^{\prime}(-\frac{1}{2})(t_{2}+\frac{1}{2})^{2}$. Since $T_{1}(t)=t$ and $T_{2}(t)=2t^{2}-1$, it easily follows that $\tilde{g}$ is CPSD if and only if $b_{i,j}\geq 0$ for $(i,j)\neq 0$. From (65), it follows that $b_{1,2}=-\frac{1}{2}\tilde{f_{1}}(-1)\tilde{f_{2}}[-1,\frac{1}{2},\frac{1}{2}]+c$. Observing that $q\geq 0$ on $1,1]^{2}$, we choose $c=\frac{1}{2}\tilde{f_{1}}(-1)\tilde{f_{2}}[-1,\frac{1}{2},\frac{1}{2}]$ as small as possible in which case $b_{1,2}=0$.
@@
The next lemma which will be used to prove $a_{1,0}>0$ as well as being a first step in establishing that $\tilde{g}\leq\tilde{F}$ on $1,1]^{2}$.
@@
5.1 $\tilde{F}\geq\tilde{g}$ on $1,1]\times([-1,-1/2]\cup[1/2,1])$
The following lemma, proved in the appendix, establishes several necessary inequality conditions for $\tilde{g}\leq\tilde{F}$. We next show the inequality holds on $1,1]\times([-1,-1/2]\cup[1/2,1])$.